# Formatting Instructions For NeurIPS 2025

### v1

@@ Abstract
Video large language models (video LLMs) excel at video comprehension but face significant computational inefficiency due to redundant video tokens. Existing token pruning methods offer solutions. However, approaches operating within the LLM (inner-LLM pruning), such as FastV, incur intrinsic computational overhead in shallow layers. In contrast, methods performing token pruning before the LLM (outer-LLM pruning) primarily address spatial redundancy within individual frames or limited temporal windows, neglecting the crucial global temporal dynamics and correlations across longer video sequences. This leads to sub-optimal spatio-temporal reduction and does not leverage video compressibility fully. Crucially, the synergistic potential and mutual influence of integrating these strategies remain unexplored. To further reduce redundancy, we introduce HoliTom, a novel training-free holistic token merging framework. HoliTom employs outer-LLM pruning through global redundancy-aware temporal segmentation, followed by spatial-temporal merging to reduce visual tokens by over 90%, significantly alleviating the LLM’s computational burden. Complementing this, we introduce a robust inner-LLM token similarity-based merging approach, designed for superior performance and compatibility with outer-LLM pruning. Evaluations demonstrate our method’s promising efficiency-performance trade-off on LLaVA-OneVision-7B, reducing computational costs to 6.9% of FLOPs while maintaining 99.1% of the original performance. Furthermore, we achieve a 2.28$\times$ reduction in Time-To-First-Token (TTFT) and a 1.32$\times$ acceleration in decoding throughput, highlighting the practical benefits of our integrated pruning approach for efficient video LLMs inference.
@@ 1 Introduction
Video large language models (video LLMs) [19, 53, 41, 7, 8, 21, 23, 46, 50] have shown remarkable potential in understanding complex video content. However, their practical deployment is hindered by significant computational inefficiency. This inefficiency stems from processing high volumes of video tokens generated by encoding sampled frames, leading to substantial overhead, particularly due to the quadratic complexity of the attention mechanism in the LLMs. For videos with numerous frames, the input token count can easily reach tens of thousands, making inference computationally expensive. While prior works [6, 48, 45, 36] have explored model compression and token pruning, achieving a desirable balance between efficiency and performance in video tasks remains an open challenge. Thus, developing effective methods to reduce video token redundancy while preserving critical semantic information is crucial for the widespread adoption of video LLMs.
@@
Empirical evaluations validate the effectiveness of our proposed method in achieving a compelling efficiency-performance trade-off. Specifically, as shown in Fig. 1 (right), on the LLaVA-OneVision-7B model [19], our approach reduces computational costs to just 6.9% of the original FLOPs while remarkably preserving 99.1% of the original model’s performance. Moreover, we observe significant gains in inference efficiency, achieving a 2.28$\times$ reduction in Time-To-First-Token (TTFT) and a 1.32$\times$ acceleration in decoding throughput. These results clearly demonstrate the substantial practical advantages of our holistic token merging framework for efficient video LLM inference.
@@ 2.1 Video Large Language Models
The rapid progress of multimodal large language models has led to the integration of video encoders, creating video LLMs that excel in video understanding and question answering tasks [46, 14, 19, 2, 41, 3, 21, 22, 23, 50, 35, 37, 17, 1]. However, the substantial number of tokens generated by processing numerous video frames hinders inference efficiency, thereby impeding the widespread adoption of video LLMs. Existing approaches have attempted to mitigate this issue. For instance, VideoLLaMA [50] employs a Q-Former module [20] to aggregate video tokens, while MovieChat [35] introduces a memory module to merge and store token representations. Although pooling mechanisms in LLaVA-OneVision [19] reduce token counts, each video frame still produces hundreds of tokens for downstream processing. Consequently, handling tens of thousands of visual tokens for long video inputs substantially increases inference time and memory consumption. While works such as VILA [25] and NVILA [26] aim to optimize token usage, these methods often require model fine-tuning, demanding considerable hardware resources [16, 23, 25, 26, 40]. This underscores a critical need for developing more efficient, training-free token compression methods specifically for video LLMs, bypassing the need for costly model adaptations and significant hardware investment.
@@ 3.2 Global Redundancy-Aware Temporal Merging
For a temporal segment defined by a start frame $t_{s}$ and an end frame $t_{e}$ (covering $[t_{s},t_{e})$), the total number of prunable tokens, $g(t_{s},t_{e})$, is calculated. This involves counting tokens $N(t_{s},t_{e})$ that are consecutively redundant across all frames from $t_{s}$ to $t_{e}-1$, and then multiplying by the number of subsequent frames $(t_{e}-t_{s}-1)$ within the segment where these tokens can be pruned. Our method prunes the redundant by merging these subsequent occurrence tokens into their first appearance at start frame $t_{s}$, treating them as temporal redundant tokens, as shown in Fig. 2. The formulation is:
@@
Given a video of $B$ frames, our objective is to find a segmentation into $K$ consecutive segments $[t_{i},t_{i+1})$ (with $t_{1}=1$, $t_{K+1}=B+1$, and $t_{i}<t_{i+1}$) that maximizes the total prunable features:
@@
This optimization is solved using dynamic programming to achieve global optimization. Let $dp[i]$ be the maximum prunable features for a video ending at frame $i$ (exclusive, i.e., considering frames $11,...,i-$ 1), where frame $i$ marks the exclusive end of the last segment. The value $prev[i]$ stores the optimal starting frame $j^{*}$ of this final segment $[j^{*},i)$. The state transition is given by:
@@
The base case is $dp[1]=0$. The maximum prunable features for the entire video are $dp[B+1]$. The optimal segmentation is reconstructed by backtracking from $B+1$ using the $prev$ array.
@@ 3.3 Spatial Merging
where $d$ is the state dimension. Token importance is quantified by averaging the attention weights each token receives from all other tokens within the same frame in the vision tower, yielding an score vector $A_{\text{avg}}\in\mathbb{R}^{B\times N_{v}}$. Tokens receiving higher average attention are considered more salient.
@@
The computation of vision tower attention is intra-frame. Averaging these attention weights across frames lacks theoretical justification, invalidating the attention-based method for redundant temporal tokens. To process these features, we employ a cluster-based merging method utilizing density peak clustering based on $k$-nearest neighbors (DPC-KNN) [12, 31]. Given a set of $N$ redundant temporal tokens $[v_{1},v_{2},...,v_{N}]$ within the first frame of the segmentation. For each token $v_{i}$, we calculate its local density $\rho_{i}$, distance to the closest higher-density token $\delta_{i}$ and the final density score $\gamma_{i}=\rho_{i}\times\delta_{i}$:
@@ 3.4 Inner-LLM Merging
Inefficient visual attention in large vision language models has been widely discussed [6, 45]. Existing methods, such as FastV [6] and PDrop [45], directly discard redundant visual tokens, which may lead to performance degradation due to information loss. Unlike these approaches, our proposed method addresses it by merging the information from potentially redundant tokens instead of simply discarding them. Specifically, at the K-th layer of the LLM, to reduce the number of visual tokens by R%, we employ a token selection strategy based on attention scores. We use the attention weights of the last token to rank all vision tokens at layer K. The Rof visual tokens exhibiting the lowest attention scores are identified as candidates for merging. We find its most similar visual token within the set of tokens designated for retention. For a retained token $v_{r}$ and its associated set of low attention tokens $V_{m}=\{v_{m_{1}},v_{m_{2}},...,v_{m_{n}}\}$. The updated retained token $v_{r}^{\prime}$ is:
@@ 4.1 Experimental Settings
Compared Methods. We compare our proposed HoliTom against 6 strong training-free baselines: 1) FastV [6], identifies key tokens during prefilling using attention scores between predicted and vision tokens; 2) PDrop [45], prunes visual tokens within partitioned LLM stages, guided by image and instruction tokens; 3) Visionzip [48], prunes tokens before LLM via spatial token merging; 4) DyCoke [36], employs temporal merging before LLM and dynamic KV cache pruning in decoding; 5) PruneVid [15], minimizes video redundancy via spatial-temporal token clustering; and 6) FastVID [33], a concurrent work, partitions videos and applies density-based token pruning. Due to the lack of public code, we compare FastVID to its reported results. For all other baselines and our method, experiments use their open-source code under identical hardware condition.
@@
Method Prefilling FLOPs (T) $\downarrow$ FLOPs Ratio $\downarrow$ Before LLM Retained Ratio MVBench $\uparrow$ EgoSchema $\uparrow$ LongVideo Bench $\uparrow$ VideoMME $\uparrow$ Avg. $\uparrow$
@@ 4.2 Main Results
Model Method Prefilling FLOPs (T) $\downarrow$ FLOPs Ratio $\downarrow$ Before LLM Retained Ratio MVBench $\uparrow$ EgoSchema $\uparrow$ LongVideo Bench $\uparrow$ VideoMME $\uparrow$ Avg. $\uparrow$
@@
LLaVA- OneVision-72B Vanilla 429.3 100% 100% 60.9 61.1 62.7 65.7 62.6 100
@@
LLaVA- Video-7B Vanilla 80.2 100% 100% 60.4 57.2 58.9 64.3 60.2 100
@@
Figure 5: Achieving superior inference. "Other" indicates token pre-processing time (e.g., pooling). Our proposed method reduces Time-To-First-Token (TTFT) by 2.28$\times$ and achieves 1.32$\times$ higher decoding throughput, outperforming all other token compression methods and the vanilla model.
@@ 4.3 Efficiency Results
Fig. 5 summarizes the impact of various token compression methods on the inference efficiency of video LLMs. As shown, all the evaluated methods demonstrably reduce LLM prefilling time. Our method, in particular, reduces it to just 13.7% of the original. For VisionZip [48], PruneVid [15], and our HoliTom require token pre-processing, which introduces additional "other" time. Furthermore, both PruneVid and our method produce a variable number of tokens per frame, complicating batch processing, which contributes to extra overhead. Our method, designed to maximize temporal redundancy pruning, leads to finer-grained segmentation, further influencing this observation. Nevertheless, our method achieves the maximum reduction in Time-To-First-Token latency, reducing it by 2.28$\times$, while maintaining optimal performance. Although we did not specifically optimize for decoding, our model’s decoding speed still benefits from the reduced number of vision tokens. Our throughput increased by 1.32$\times$ compared to the original model, the highest among all methods evaluated.
@@ 4.4 Ablation Study
Method Prefilling FLOPs (T) $\downarrow$ FLOPs Ratio $\downarrow$ Before LLM Retained Ratio MVBench $\uparrow$ EgoSchema $\uparrow$ LongVideo Bench $\uparrow$ VideoMME $\uparrow$ Avg. $\uparrow$
@@
Methods MVBench EgoSchema LongVideo Bench VideoMME Avg.
@@ 5 Conclusion
This paper presents HoliTom, a new training-free holistic token merging framework for boosting the efficiency of video LLMs by effectively handling redundant visual tokens. HoliTom achieves this through a synergistic integration of outer-LLM spatio-temporal reduction, drastically reducing initial token counts, and a robust inner-LLM token merging mechanism tailored for compatibility and further optimization. Evaluated on prominent video LLMs, HoliTom achieves a state-of-the-art efficiency-performance trade-off, substantially reducing computational costs (e.g., to 6.9% FLOPs) while preserving high performance (e.g., 99.1% accuracy), and accelerating inference (2.28$\times$ TTFT, 1.32$\times$ throughput). These results underscore the effectiveness of HoliTom in enabling practical and efficient video LLMs inference for complex, long-form video understanding.
@@ Appendix A Supplemental Implementation Details
Our method is implemented on the LLaVA-OneVision-7B/72B [19] and LLaVA-Video-7B [53] models. Evaluation utilized NVIDIA A100 (80GB) GPUs; inference was performed on an NVIDIA RTX A6000 GPU. To ensure a fair comparison of computational cost, we used total prefilling FLOPs as the primary metric. Baselines are configured for comparable FLOPs: FastV [6] prunes 80% of tokens at layer 2; PDrop [45] retains 50%, 25%, and 12.5% of vision tokens at layers 2, 7, and 14, respectively; VisionZip [48] and PruneVid [15] maintain a consistent proportion of input tokens with our method. Performance results for FastVID [33] are adopted directly from their original paper. For our proposed method, the default threshold $\tau$ is 0.8. In experiments targeting a 10% compression ratio, $\tau$ was set to 0.65. Experimental setups include pruning K=18 layers of the 7B model and K=60 layers of the 72B model, both at a ratio of R=50%. Following the official LLaVA-OneVision specifications, the default input video frames are 32 and $N_{v}=196$. For LLaVA-Video, the default input consisted of 64 video frames with $N_{v}=169$. All benchmark evaluations are performed using the LMMs-Eval [51, 18].

### v3

# HoliTom [Uncaptioned image] : Holistic Token Merging for Fast Video Large Language Models

@@ Abstract
[…]
@@ 1 Introduction
Video large language models (video LLMs) [21, 62, 48, 7, 8, 23, 26, 55, 59, 15, 43] have shown remarkable potential in understanding complex video content. However, their practical deployment is hindered by significant computational inefficiency. This inefficiency stems from processing high volumes of video tokens generated by encoding sampled frames, leading to substantial overhead, particularly due to the quadratic complexity of the attention mechanism in the LLMs. For videos with numerous frames, the input token count can easily reach tens of thousands, making inference computationally expensive. While prior works [6, 57, 53, 41, 30] have explored model compression and token pruning, achieving a desirable balance between efficiency and performance in video tasks remains an open challenge. Thus, developing effective methods to reduce video token redundancy while preserving critical semantic information is crucial for the widespread adoption of video LLMs.
@@
[…]
@@ 2.1 Video Large Language Models
The rapid progress of multimodal large language models has led to the integration of video encoders, creating video LLMs that excel in video understanding and question answering tasks [55, 14, 21, 2, 48, 3, 23, 24, 26, 59, 40, 42, 19, 1, 39, 25]. However, the substantial number of tokens generated by processing numerous video frames hinders inference efficiency, thereby impeding the widespread adoption of video LLMs. Existing approaches have attempted to mitigate this issue. For instance, VideoLLaMA [59] employs a Q-Former module [22] to aggregate video tokens, while MovieChat [40] introduces a memory module to merge and store token representations. Although pooling mechanisms in LLaVA-OneVision [21] reduce token counts, each video frame still produces hundreds of tokens for downstream processing. Consequently, handling tens of thousands of visual tokens for long video inputs substantially increases inference time and memory consumption. While works such as VILA [28] and NVILA [29] aim to optimize token usage, these methods often require model fine-tuning, demanding considerable hardware resources [18, 26, 28, 29, 47]. This underscores a critical need for developing more efficient, training-free token compression methods specifically for video LLMs, bypassing the need for costly model adaptations and significant hardware investment.
@@ 3.2 Global Redundancy-Aware Temporal Merging
For a temporal segment defined by a start frame $t_{s}$ and an end frame $t_{e}$ (covering $t_{s},t_{e})$), the total number of prunable tokens, $g(t_{s},t_{e})$, is calculated. This involves counting tokens $N(t_{s},t_{e})$ that are consecutively redundant across all frames from $t_{s}$ to $t_{e}-1$, and then multiplying by the number of subsequent frames $(t_{e}-t_{s}-1)$ within the segment where these tokens can be pruned. Our method prunes the redundant by merging these subsequent occurrence tokens into their first appearance at start frame $t_{s}$, treating them as temporal redundant tokens, as shown in Fig. 2. The formulation is:
@@
Given a video of $B$ frames, our objective is to find a segmentation into $K$ consecutive segments $t_{i},t_{i+1})$ (with $t_{1}=1$, $t_{K+1}=B+1$, and $t_{i}<t_{i+1}$) that maximizes the total prunable features:
@@
This optimization is solved using dynamic programming to achieve global optimization. Let $dp[i]$ be the maximum prunable features for a video ending at frame $i$ (exclusive, i.e., considering frames 1,,i11,...,i-1), where frame $i$ marks the exclusive end of the last segment. The value $prev[i]$ stores the optimal starting frame $j^{*}$ of this final segment $j^{*},i)$. The state transition is given by:
@@
The base case is $dp[1]=0$. The maximum prunable features for the entire video are $dp[B+1]$. The optimal segmentation is reconstructed by backtracking from $B+1$ using the prevprev array.
@@ 3.3 Spatial Merging
where $d$ is the state dimension. Token importance is quantified by averaging the attention weights each token receives from all other tokens within the same frame in the vision tower, yielding a score vector $A_{\text{avg}}\in\mathbb{R}^{B\times N_{v}}$. Tokens receiving higher average attention are considered more salient.
@@
The computation of vision tower attention is intra-frame. Averaging these attention weights across frames lacks theoretical justification, invalidating the attention-based method for redundant temporal tokens. To process these features, we employ a cluster-based merging method utilizing density peak clustering based on $k$-nearest neighbors (DPC-KNN) [12, 35]. Given a set of $N$ redundant temporal tokens $v_{1},v_{2},...,v_{N}]$ within the first frame of the segmentation. For each token $v_{i}$, we calculate its local density $\rho_{i}$, distance to the closest higher-density token $\delta_{i}$ and the final density score $\gamma_{i}=\rho_{i}\times\delta_{i}$:
@@ 3.4 Inner-LLM Merging
Inefficient visual attention in large vision language models has been widely discussed [6, 53]. Existing methods, such as FastV [6] and PDrop [53], directly discard redundant visual tokens, which may lead to performance degradation due to information loss. Unlike these approaches, our proposed method addresses it by merging the information from potentially redundant tokens instead of simply discarding them. Specifically, at the K-th layer of the LLM, to reduce the number of visual tokens by R%, we employ a token selection strategy based on attention scores. We use the attention weights of the last token to rank all vision tokens at layer K. The R% of visual tokens exhibiting the lowest attention scores are identified as candidates for merging. We find its most similar visual token within the set of tokens designated for retention. For a retained token $v_{r}$ and its associated set of low attention tokens $V_{m}=\{v_{m_{1}},v_{m_{2}},...,v_{m_{n}}\}$. The updated retained token $v_{r}^{\prime}$ is:
@@ 4.1 Experimental Settings
Compared Methods. We compare our proposed HoliTom against 6 strong training-free baselines: 1) FastV [6], identifies key tokens during prefilling using attention scores between predicted and vision tokens; 2) PDrop [53], prunes visual tokens within partitioned LLM stages, guided by image and instruction tokens; 3) Visionzip [57], prunes tokens before LLM via spatial token merging; 4) DyCoke [41], employs temporal merging before LLM and dynamic KV cache pruning in decoding; 5) PruneVid [17], minimizes video redundancy via spatio-temporal token clustering; and 6) FastVID [37], a concurrent work, partitions videos and applies density-based token pruning. Due to the lack of public code, we compare FastVID to its reported results. For all other baselines and our method, experiments use their open-source code under identical hardware condition.
@@
Method
Prefilling
FLOPs (T) $\downarrow$
FLOPs
Ratio $\downarrow$
Before LLM
Retained Ratio
MVBench
$\uparrow$
EgoSchema
$\uparrow$
LongVideo
Bench $\uparrow$
VideoMME
$\uparrow$
Avg. $\uparrow$
@@ 4.2 Main Results
Model Method
Prefilling
FLOPs (T) $\downarrow$
FLOPs
Ratio $\downarrow$
Before LLM
Retained Ratio
MVBench
$\uparrow$
EgoSchema
$\uparrow$
LongVideo
Bench $\uparrow$
VideoMME
$\uparrow$
Avg. $\uparrow$
@@
LLaVA-
OneVision-72B
Vanilla 429.3 100% 100% 60.9 61.1 62.7 65.7 62.6 100
@@
LLaVA-
Video-7B
Vanilla 80.2 100% 100% 60.4 57.2 58.9 64.3 60.2 100
@@
[…]
@@ 4.3 Efficiency Results
[…]
@@ 4.4 Ablation Study
Method
Prefilling
FLOPs (T) $\downarrow$
FLOPs
Ratio $\downarrow$
Before LLM
Retained Ratio
MVBench
$\uparrow$
EgoSchema
$\uparrow$
LongVideo
Bench $\uparrow$
VideoMME
$\uparrow$
Avg. $\uparrow$
@@
Methods MVBench EgoSchema
LongVideo
Bench
VideoMME Avg.
@@ 5 Conclusion
[…]
Acknowledgment
This paper is supported by the Young Scientists Fund of the National Natural Science Foundation of China (No. 62506305), Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (No. 2024R01007), Key Research and Development Program of Zhejiang Province (No. 2025C01026), and Scientific Research Project of Westlake University (No. WU2025WF003).
@@ Appendix A Supplemental Implementation Details
Our method is implemented on the LLaVA-OneVision-7B/72B [21] and LLaVA-Video-7B [62] models. Evaluation utilized NVIDIA A100 (80GB) GPUs; inference was performed on an NVIDIA RTX A6000 GPU. To ensure a fair comparison of computational cost, we used total prefilling FLOPs as the primary metric. Baselines are configured for comparable FLOPs: FastV [6] prunes 80% of tokens at layer 2; PDrop [53] retains 50%, 25%, and 12.5% of vision tokens at layers 2, 7, and 14, respectively; VisionZip [57] and PruneVid [17] maintain a consistent proportion of input tokens with our method. Performance results for FastVID [37] are adopted directly from their original paper. For our proposed method, the default threshold $\tau$ is 0.8. In the specific experiments conducted on Qwen2.5-VL [3] with a maximum sampling of 768 frames, a lower threshold of $\tau=0.2$ was used. This adjustment accounts for the higher temporal redundancy present when sampling is dense. In experiments targeting a 10% compression ratio, $\tau$ was set to 0.65. Experimental setups include pruning K=18 layers of the 7B model and K=60 layers of the 72B model, both at a ratio of R=50%. Following the official LLaVA-OneVision specifications, the default input video frames are 32 and $N_{v}=196$. For LLaVA-Video, the default input consisted of 64 video frames with $N_{v}=169$. All benchmark evaluations are performed using the LMMs-Eval [60, 20].
Appendix B Supplemental Experimental Results
B.1 Experiments on Qwen2.5-VL with High Frame Sampling
Table 5: Comparison of state-of-the-art methods on Qwen2.5-VL 7B. Qwen2.5-VL accepts video input at 2 fps, capped at a maximum of 768 frames, maximum video token limit to 24,576 (as the technical report recommends). A/B in the # Token column indicates that A tokens are first provided to the LLM, and then compressed to B tokens during the LLM forward pass. The token number is derived from the average video token count per task within VideoMME. FastV performs full attention matrix calculation in memory, causing OOM errors due to the large number of video tokens.
Method # Token $\downarrow$ TFLOPs $\downarrow$ VideoMME $\uparrow$
Short Medium Long Overall
Qwen2.5-VL-7B 18442 100% 377.2 100% 77.4 68.1 55.6 67.0 100%
FastV [6] 18442 / 9221 100% / 50.0% 170.4 45.2% OOM
Visionzip [57] 9221 50.0% 154.5 41.0% 74.9 66.6 55.7 65.7 98.1%
PruneVid [17] 9173 49.7% 153.5 40.7% 72.3 64.8 54.7 63.9 95.4%
HoliTom (w/o M) 6513 34.9% 102.0 27.0% 74.4 66.4 56.4 65.8 98.2%
FastV [6] 18442 / 4610 100% / 25.0% 90.7 24.0% OOM
Visionzip [57] 4610 25.0% 68.7 18.2% 73.1 63.3 55.9 64.1 95.7%
PruneVid [17] 4632 25.1% 69.1 18.3% 69.3 61.1 53.2 61.2 91.3%
HoliTom (w/o M) 4504 24.4% 66.9 17.7% 72.7 65.7 56.1 64.8 96.7%
Existing models like LLaVA-OV [21] and LLaVA-Video [62] utilize a fixed input of 32/64 video frames, each resized to a static resolution (Tab. 23). In contrast, frontier models, such as Qwen2.5-VL [3], introduce advanced features including FPS frame sampling, which extend input sequences (up to 768 frames), and dynamic resolution support. These new capabilities pose new challenges to existing token compression methods in maintaining performance for video understanding tasks.
As shown in Tab. 5, HoliTom surpasses state-of-the-art methods across both token compression rates, especially for long videos. Due to Out-of-Memory (OOM) issues arising from the full attention matrix calculation, we were unable to report results for FastV and the inner-LLM merging execution.
Figure 9: Fine-grained object tasks. Performance on MVBench object subtasks (object existence (OE), object interaction (OI), and object shuffle (OS)) improves with more aggressive token compression.
Method FLOPs (%) MVBench
OE OI OS Overall
OV 7B 100 57.5 84.0 35.5 58.3
+HoliTom 17.4 61.0 83.5 36.5 58.4
+HoliTom 14.2 63.0 84.0 37.5 58.7
+HoliTom 10.5 60.5 84.5 38.0 58.1
Figure 10: Improved performance with enhanced efficiency. Using more input frames, token compression boosts performance while controlling computational overhead.
Method # Frame FLOPs (%) Avg. Score
OV 7B 32 100 58.4
+HoliTom 32 12.7 58.5
+HoliTom 64 26.5 58.9
+HoliTom 128 56.6 59.3
B.2 Impact of Token Compression on Fine-Grained Object Understanding
HoliTom applies aggressive video token compression. Does this aggressive compression impair the model’s ability to comprehend fine-grained details? Tab. 10 presents the performance results on selected subtasks from the MVBench benchmark (originally detailed in Table 2).
At compression rates ranging from 10% to 25%, HoliTom demonstrates increases in performance. This result would be improbable if HoliTom discards critical information about small objects. Instead, this outcome provides evidence of HoliTom’s robust ability to retain fine-grained details.
B.3 Enhanced Performance with Reduced Overhead
Tab. 10 extends the findings presented in Fig. 4. By sampling more frames with HoliTom while maintaining constant or even reduced total FLOPs, we achieve better performance compared to a vanilla model operating on fewer frames. We also observe that as the number of input frames increases, the computational overhead contributed by the vision encoder becomes a non-negligible factor. This performance and efficiency trade-off is further illustrated in Fig. 5 of our paper.